OpenStudy (anonymous):

The Look Alike Doll company makes dolls resemble people. Pam's mom is having a doll made to look like pam. Pam is 5 feet tall. The doll will be 2 1/2 feet tall. what is the ratio of the doll's eheight to pam's height? a. 5 ft : 2 1/2 ft b. 2ft to 5 ft c. 1/5 ft d. 1/2 ft

OpenStudy (dumbcow):

ratio = 2.5/5 = 1/2

OpenStudy (anonymous):

thx! (yay i got it right)